OpenStudy (anonymous):

Solve. Round your answer to the nearest hundredth. x^2 + 45 = 0

OpenStudy (campbell_st):

slight problem here as this question will have a complex solution

OpenStudy (anonymous):

\[x=\pm \sqrt{-45}\] which comes out pretty horribly... \[x=\pm 6.708i\]

OpenStudy (campbell_st):

\[x = \pm i \sqrt{45}\]

Directrix (directrix):

x^2 + 45 = 0 x^2 = -45 x = √ (-45) or x = -√ (-45) x = √(-1) * √45 or x = -√(-1) * √45 x = i √45 or x = -i √45 x = 6.71 i or - 6.71 i